OpenStudy (anonymous):

(12x^2yz)^2 ?? how should i solve this?

OpenStudy (campbell_st):

well its really \[12^2 \times (x^2)^2 \times y^2 \times z^2\] simplify the power of x and find 12^2 for the answer

OpenStudy (campbell_st):

the power of 2 operates on each part of the term

OpenStudy (anonymous):

So would it be 144x^4 y^2 z^2?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

what if it was (u^3 over 6) ^2?

OpenStudy (campbell_st):

well then its \[\frac{(u^3)^2}{6^2}\]

OpenStudy (campbell_st):

you just need to simplify it

OpenStudy (anonymous):

what would u's power be, 6 or 5?

OpenStudy (campbell_st):

the law for power of a power is \[(x^a)^b = x^{a \times b}\] this is the law you need to simplify the power